OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (0, 6) and (1, 3)?

OpenStudy (anonymous):

(y - 3) = -3(x - 1) (y + 3) = 3(x + 1) (y + 3) = -3(x + 1) (y - 3) = 3(x + 6)

OpenStudy (anonymous):

You don't have to tell me the answer, just help me solve it!

OpenStudy (zale101):

use the slope formula \[slope=\frac{ y_2-y_1 }{ x_2-x_1 }\] after you find the slope (m) use it to thepoint slope formula \[y-y_2=m(x-x_1)\]

OpenStudy (anonymous):

thanks! You're so helpful!

OpenStudy (zale101):

of course we'll walk you through :) this is what's openstudy for

OpenStudy (anonymous):

I think I got it now, it's (y - 3) = -3(x - 1)

OpenStudy (anonymous):

I appreciate your help though thank you :)

OpenStudy (zale101):